# Position IDs Matter: An Enhanced Position Layout for Efficient Context Compression in Large Language Models

### v3

@@ 4.3 Pretraining Results
Through pretraining LLMs have learned to compress context into memory tokens, allowing evaluation over memory tokens for its reconstruction and language modeling capability. The evaluation methodology is widely adopted for soft prompting Ge et al. (2024); Li et al. (2024a) and we expect EPL to bring a similar improvement to the fine-tuning settings since EPL incorporates useful prior to reconstruction and language modeling through its position layouts. The reconstruction quality and language modeling capability are evaluated using BLEU-4 Papineni et al. (2002) and perplexity (PPL), respectively [^9] Reconstruction texts are generated via greedy search..
@@ 4.4 Applications to Multimodal Models
As EPL can be applied to all applications that compress context into special tokens, in this subsection, we showcase its application in multimodaility. We follow VoCo-LLaMA Ye et al. (2025) for visual question-answering tasks. Given a triplet $(I,Q,A)$, the model encodes image $I$ into a sequence of 576 visual tokens $V^{t}=\{vt_{0},vt_{1},...,vt_{575}\}$, and subsequently compresses $V^{t}$ into the KV values of Vision Compression (VoCo) tokens. VoCo-LLaMA adopts a single training stage akin to fine-tuning stage in 500xCompressor and employs a single-forward via an attention mask (see Figure 8 for the attention mask detail) to prevent $Q$ and $A$ from directly accessing $V^{t}$. VoCo-LLaMA uses DPL and we follow its experimental setup to examine the effect of changing DPL to EPL. The position layout changes are illustrated on top of Figure 8.
We evaluate VoCo-LLaMA with 128 VoCo tokens (i.e. 4.5x compression ratio) and report performance on multimodal benchmarks. As shown in Table 5, VoCo-LLaMA combined with EPL significantly outperforms both its DPL counterpart from our reproduction and the results reported by Ye et al. (2025). We observe improvement across all three evaluated tasks, validating again the universality of EPL.
@@ 4.5 Ablation Studies
Table 6: Ablation study on EPL integration. +UPL applies only UPL to encoder; UPL & CPL (i.e. EPL) applies both UPL and CPL. Reported values are averages. More ablation results in Tables 18 and 19.
We conduct ablation studies to analyze the independent effects of UPL and CPL in EPL (see Section 3.2 for detailed description of UPL and CPL). LM perplexity and AE BLEU are measured under the same experimental settings as section 4.3 while out-of-domain performance is measured same as section 4.2. We show the results in Table 6 which confirm that:
@@
For all testing cases (AE and LM during pretraining and MRQA out-of-domain performance for fine-tuning) over all architectures (ICAE and 500xCompressor), UPL improves performance compared to its counterparts. We note that the improvement is more significant for the weaker ICAE model. For example, the MRQA out-of-domain ROUGE-1 F1 improves 3.72 while the improvement was 1.07 for 500xCompressor. The results confirm that by carefully designing the memory position layout for compression tasks, the memory tokens obtain useful compression inductive prior, improves the compression efficiency (i.e. AE task performance) and consequently downstream task performance as shown by MRQA out-of-domain performance.
@@
[…]
@@ 4.6 Scalability Results
We conduct experiments at 3B and 8B scales from the same model family to verify method scalability. Table 7 shows that EPL achieves performance improvements at all scales and the improvement does not attenuate with larger scale. For example, at 8B scale, the MRQA ROUGE-1 F1 increased 14.03 for ICAE and increased 1.89 for 500xCompressor, which are similar to the improvement observed in 1B case in Table 4, holding promise for the method’s effectiveness on large-scale language models. More detailed results on MRQA can be found in Table 16 and 17.
@@ 5.2 Attention Visualization
To verify whether the performance improvement we observe previously is indeed due to more suitable attention patterns for compression tasks idealized in Figure 1, we visualize the summed attention matrices of all attention heads in the second and final layers of the ICAE-3B model (Figure 6 and Figure 7) after the model has been fine-tuned on MRQA tasks.
For the second layer, we observe that the memory token under UPL attends to its surrounding [^10] In terms of position IDs, not physical token positions. context tokens, forming a slope (central top of Figure 6), contrast to DPL that only exhibits self-attention among memory tokens, showing that the attention adheres to our specified prior through EPL even after learning [^11] The angle of the slope is around $11^{\circ}$ ($\arctan(1/\textbf{5})=11.3^{\circ}$ at a compression rate of 5), which is consistent with our UPL conception illustrated in Figure 1. . In the final layer (Figure 7), UPL maintains the slope pattern, while DPL overcomes the position encoding resistance and attends to distant context tokens through learning without a clear pattern.
@@ 5.3 Learnable PEs
Our results rely on local bias of PEs that are explicitly defined in  Vaswani et al. (2017); Su et al. (2023); Press et al. (2022). We examine in this section empirically the properties of learnable position encodings Devlin et al. (2019); Raffel et al. (2020).
Figure 9 shows the cosine similarity of BERT’s Devlin et al. (2019) position embeddings where we observe significantly higher similarity between adjacent position embeddings compared to distant tokens. Similarly, The bias values of T5 Bias Raffel et al. (2020) decay with increasing relative distance as shown in Figure 2, indicating that attention is stronger amongst close tokens.
The BERT’s [CLS] token (position ID 0) shows a special pattern as shown in Figure 10: contrast to other tokens, its cosine similarities are not higher with its adjacent tokens. Given [CLS] token behaves like a compressed memory token trained using next sentence prediction, the phenomenon motivates our current work, suggesting that different priors should be given to special tokens for best performance.
@@ 6 Related Work
GIST Mu et al. (2023) trained LLMs with modified attention mechanisms (similar to Figure 8) to compress prompt information into a few gist tokens. AutoCompressor Chevalier et al. (2023) trained an LLM to recursively compress long prompts by combining compressed tokens with new sub-prompts in each iteration, ultimately collecting all compressed tokens to form a compact representation. ICAE Ge et al. (2024) introduced an AE task enabling LLMs to pre-train compression capabilities on large-scale corpora, requiring only minimal parameter tuning for the encoder while freezing the decoder. 500xCompressor Li et al. (2024b) built upon ICAE by changing the information carrier from memory token outputs to memory tokens’ KV values. UniICL Gao et al. (2024b) and SelfCP Gao et al. (2024a) freeze both the encoder and decoder, training only a connector module to transform the encoder’s output memory tokens into decoder inputs. VoCo-LLaMA Ye et al. (2025) is the first to use LLMs for compressing visual tokens. Similar to GIST, it compresses visual token information into VoCo tokens through modified attention masks, outperforming methods like Q-Former Li et al. (2023) and average pooling with linear projection Li et al. (2024a). None of these methods discussed the impact of position layout and our EPL can be applied to all these soft prompt methods.
@@
Although we are not aware of position layout work in the compression domain, we find related work in multimodality by considering memory tokens as another modality. Due to images being two-dimensional and text being one-dimensional, handling mixed image-text positional layout remains an open question. Current mainstream approaches in multimodality flatten 2D images into 1D sequences Liu et al. (2023b, a); Bavishi et al. (2023); Lu et al. (2024); Sun et al. (2024). Seeking a more elegant solution, Su (2024) proposed RoPE-Tie by placing visual and text tokens along the diagonal $(y=x)$ in 2D space. Although Su (2024) does not thoroughly validate their design, as the approach maintains the sequence order between modalities and the internal locality of images and text respectively, our empirical results suggest that the design can results in better performance. Qwen2-VL Wang et al. (2024) adopt similar designs in video domains.
@@ 7 Conclusion
We examine position layout, an understudied topic in context compression, and propose EPL for soft prompt methods. Our proposed position layout improves over default position layout by bringing memory tokens close to its context and at the same time maintains the logical sequence token ordering amongst context, memory tokens and the subsequent tokens. Extensive experiments demonstrate that EPL improves the compression efficiency and downstream task performance over different architectures across different modalities.
Given LLM’s ubiquitous usage, we believe that carefully examining position layout will give fruits to problems beyond context compression. We hope the success of EPL fosters research in this under explored area.
8 Limitations
Throughout our experiments, we have tested and confirmed EPL effectiveness under 5x compression ratio for text reconstruction and QA tasks and 4.5x compression ratio for visual QA. There remains question if our method is still effective at high compression ratio. We don’t have a positive outlook on this question.
First, EPL, requiring memory tokens to achieve uniform coverage across the context, intrinsically aligns with lossless compression scenarios (i.e., autoencoding tasks) which assumes that “all tokens in the context are equally important”; however, under high compression ratio (lossy compression), with information carrier capacity being limited, focus should be placed on important tokens, which violates our “equally important” assumption. Empirically, Figure 10 shows that the BERT’s [CLS] would have higher attention to some tokens without a clear pattern. Our preliminary results in Table 10 on VoCo-LLaMA also shows that EPL does not demonstrate any significant gains when used under high compression ratio (288x).
The analysis suggests that high ratio lossy compression may require compression mechanism that goes beyond our current work. At higher level, this suggests when adapting position layout methods to different application scenarios, the success can highly depend on whether the conceived layout captures the underlying prior characteristics of specific tasks.
@@ Appendix A Detailed Position Layout
Detailed position layouts are provided in this section. Table 11 illustrates the position layout for LLMs without compression under standard conditions, reflecting natural language priors. Table 12 details the encoder’s position layout. For the decoder, position layouts are described in Tables 13, 14, and 15, with slight variations depending on the specific task (AE, LM, or QA). Each of these tables (Tables 12, 13, 14, and 15) includes a formula at the top representing the generalized layout. Below the formula, an illustrative example is given using a context length $|C|=p=1020$, chunk size $L=510$, compression ratio $r=5$, total sequence length $|X|=2040$, question length $|Q|=50$, and answer length $|A|=5$.
@@ H.1 Sinusoidal Position Encoding
The sinusoidal position encoding Vaswani et al. (2017); Su et al. (2023) is given by:
$PE_{(pos,2i)}=\sin\left(\frac{pos}{10000^{2i/d_{model}}}\right)$
$PE_{(pos,2i+1)}=\cos\left(\frac{pos}{10000^{2i/d_{model}}}\right)$
Consider two nearby positions, $pos$ and $pos+\delta$, where $\delta$ is a small number.
@@
$\sin(x+\epsilon)\approx\sin(x)\quad\text{for small }\epsilon$
$\cos(x+\epsilon)\approx\cos(x)\quad\text{for small }\epsilon$
@@ H.2 Learnable Position Encodings
DPL 0 1 $L-1$ $L$ $L+1$ $L+|M|-1$
EPL $(i-1)L+1$ $(i-1)L+2$ $iL$ $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor(b+(|M|-1)r\right\rceil$
@@
[…]
@@
DPL(ICAE) 0 1 $k|M|-1$ $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+p$
DPL(500x) $L$ $L+1$ $L+|M|-1$ $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+p$
EPL $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor b+(k|M|-1)r\right\rceil$ 00 $1$ $2$ $p$
@@
EPL $3$ $8$ $1018$ 00 $1$ $2$ $1020$
[…]
@@
EPL $3$ $8$ $1018$ $1020$ $1021$ $1022$ $2040$
[…]
@@
EPL $3$ $8$ $1018$ $1020$ $1021$ $1070$ $1071$ $1075$
[…]
@@
Figure 6: Attention Matrix in the 2nd layer of ICAE after MRQA finetuning.

### v4

@@ 4.3 Pretraining Results
Through pretraining, LLMs have learned to compress context into memory tokens, allowing evaluation over memory tokens for their reconstruction and language modeling capability. The evaluation methodology is widely adopted for soft prompting Ge et al. (2024); Li et al. (2024a) and we expect EPL to bring a similar improvement to the fine-tuning settings since EPL incorporates useful prior to reconstruction and language modeling through its position layouts. The reconstruction quality and language modeling capability are evaluated using BLEU-4 Papineni et al. (2002) and perplexity (PPL), respectively [^9] Reconstruction texts are generated via greedy search..
@@ 4.4 Applications to Multimodal Models
As EPL can be applied to all applications that compress context into special tokens, in this subsection, we showcase its application in multimodality. We follow VoCo-LLaMA Ye et al. (2025) for visual question-answering tasks. Given a triplet $(I,Q,A)$, the model encodes image $I$ into a sequence of 576 visual tokens $V^{t}=\{vt_{0},vt_{1},...,vt_{575}\}$, and subsequently compresses $V^{t}$ into the KV values of Vision Compression (VoCo) tokens. VoCo-LLaMA adopts a single training stage akin to fine-tuning stage in 500xCompressor and employs a single-forward via an attention mask (see Figure 8 for the attention mask detail) to prevent $Q$ and $A$ from directly accessing $V^{t}$. VoCo-LLaMA uses DPL and we follow its experimental setup to examine the effect of changing DPL to EPL. The position layout changes are illustrated on top of Figure 8.
We evaluate VoCo-LLaMA with 128 VoCo tokens (i.e. 4.5x compression ratio) and report performance on multimodal benchmarks. As shown in Table 6, VoCo-LLaMA combined with EPL significantly outperforms both its DPL counterpart from our reproduction and the results reported by Ye et al. (2025). We observe improvement across all three tasks, validating the universality of EPL.
@@ 4.5 Ablation Studies
Table 7: Ablation study on EPL integration. +UPL applies only UPL to encoder; UPL & CPL (i.e. EPL) applies both UPL and CPL. Reported values are averages. More ablation results are in Tables 21 and Tables 22.
We conduct ablation studies to analyze the independent effects of UPL and CPL in EPL (see Section 3.2 for the detailed description of UPL and CPL). LM perplexity and AE BLEU are measured under the same experimental settings as in Section 4.3 while out-of-domain performance is measured in the same way as in Section 4.2. We show the results in Table 4.5 which confirm that:
@@
For all testing cases (AE and LM during pretraining and MRQA out-of-domain performance for fine-tuning) over all architectures (ICAE and 500xCompressor), UPL improves performance compared to its counterparts. We note that the improvement is more significant for the weaker ICAE model. For example, the MRQA out-of-domain ROUGE-1 F1 improves by 3.72 points while the improvement is 1.07 points for 500xCompressor. The results confirm that by carefully designing the memory position layout for compression tasks, the memory tokens obtain useful compression inductive prior, which improves the compression efficiency (i.e. AE task performance) and consequently downstream task performance as shown by MRQA out-of-domain performance.
@@
Similarly, we observe consistent improvement across the board by incorporating CPL. We note that the performance improvement for CPL on top of UPL is more prominent for the stronger 500xCompressor model; MRQA out-of-domain ROUGE-1 F1 improves by 1.20 points while the improvement is only 0.20 for ICAE. We think this is because ICAE DPL maintains the token sequential orders between $X_{\text{context}}$, [LM], and $X_{\text{completion}}$ while 500xCompressor does not. As illustrated in Table 3, while the ICAE DPL coincides with physical token sequence positions, the 500xCompressor DPL exhibits $KV$ position IDs larger than the [LM] token position ID (and some $X_{\text{completion}}$ position IDs), breaking the causal relationship between $\{X_{\text{context}},\texttt{[LM]},X_{\text{completion}}\}$ in the position layout.
@@ 4.6 Scalability Results
We conduct experiments at 3B and 8B scales from the same model family to verify method scalability. Table 4.6 shows that EPL achieves performance improvements at all scales and the improvement does not attenuate with larger scale. For example, at 8B scale, the MRQA ROUGE-1 F1 increases by 14.03 points for ICAE and by 1.89 points for 500xCompressor, which is similar to the improvement observed in 1B case in Table 4, holding promise for the method’s effectiveness on large-scale language models. More detailed results on MRQA can be found in Table 19 and 20.
@@
4.7 Result of Fine-Grained Compression
To further validate the effectiveness of EPL, we conduct additional experiments on Fine-Grained Compression (FGC) Zhang et al. (2025). FGC shares a similar motivation with our UPL, as it divides the context into finer-grained units and interleaves memory tokens among them. Since FGC places memory tokens after these fine-grained units, EPL can be naturally applied to enhance its performance by setting each memory token’s position ID to the middle position of the corresponding fine-grained unit. Given that FGC utilizes the KV cache as its compression carrier like the 500xCompressor, we incorporate FGC into the 500xCompressor architecture maintaining the same experimental setup as in Section 4.2.
LM AE Out-of-Domain
PPL BLEU F1 EM
500x-1B 11.22 93.73 45.76 31.45
+EPL 10.80 98.50 48.03 32.29
+FGC 11.03 94.49 45.52 30.40
+FGC & EPL 10.80 98.58 47.54 32.45
Table 9: Results of fine-grained compression Zhang et al. (2025).
EPL consistently enhances FGC performance
As shown in Table 9, applying EPL to the FGC yields consistent performance improvements across all evaluation metrics. These results demonstrate that EPL’s position layout adjustments provide a generalizable enhancement to existing compression methods.
4.8 Sensitivity Analysis for Compression Ratios
We conduct a sensitivity analysis of compression ratios to evaluate EPL’s robustness with experimental settings aligned with Table 4.
LM AE Out-of-Domain
PPL BLEU4 F1 EM
ICAE-1B (DPL) 5x 12.18 31.80 39.95 27.32
ICAE-1B (EPL) 5x 11.42 95.98 43.87 29.50
[2pt/2pt] ICAE-1B (DPL) 15x 12.18 8.12 34.46 23.33
ICAE-1B (EPL) 15x 12.43 52.90 36.71 23.90
[2pt/2pt] ICAE-1B (DPL) 31x 12.30 6.83 33.46 22.65
ICAE-1B (EPL) 31x 12.55 18.09 31.63 21.01
[2pt/2pt] ICAE-1B (DPL) 51x 12.18 3.81 29.99 20.39
ICAE-1B (EPL) 51x 12.55 10.19 31.46 21.01
Table 10: Results of different compression ratios.
As shown in Table 4.8, the results demonstrate that EPL significantly outperforms DPL across compression ratios up to 15x, encompassing the majority of practical application scenarios, although improvement attenuates as the compression ratio goes up. This performance range is particularly relevant, as lossless encoders typically fail to achieve compression ratios exceeding 10x Kuratov et al. (2025). EPL performance improvement becomes less consistent at higher compression ratios, which we discuss further in the Limitation Section.
@@ 5.2 Attention Visualization
To verify whether the performance improvement is indeed due to more suitable attention patterns for compression tasks idealized in Figure 1, we visualize the summed attention matrices of all attention heads in the second and final layers of the ICAE-3B model (Figure 6 and Figure 7) after the model has been fine-tuned on MRQA tasks.
For the second layer, we observe that the memory token under UPL attends to its surrounding [^10] In terms of position IDs, not physical token positions. context tokens, forming a slope (central top of Figure 6), in contrast to DPL that only exhibits self-attention among memory tokens, showing that the attention adheres to our specified prior through EPL even after learning [^11] The angle of the slope is around $11^{\circ}$ ($\arctan(1/\textbf{5})=11.3^{\circ}$ at a compression rate of 5), which is consistent with our UPL conception illustrated in Figure 1. . In the final layer (Figure 7), UPL maintains the slope pattern, while DPL overcomes the position encoding resistance and attends to distant context tokens through learning without a clear pattern.
@@ 5.3 Learnable PEs
Our results rely on the local bias of PEs that are explicitly defined in  Vaswani et al. (2017); Su et al. (2023); Press et al. (2022). We examine in this section empirically the properties of learnable position encodings Devlin et al. (2019); Raffel et al. (2020).
Figure 9 shows the cosine similarity of BERT’s Devlin et al. (2019) position embeddings where we observe significantly higher similarity between adjacent position embeddings compared to distant tokens. Similarly, the bias values of T5 Bias Raffel et al. (2020) decay with increasing relative distance as shown in Figure 2, indicating that attention is stronger among close tokens.
BERT’s [CLS] token (position ID 0) shows a special pattern as shown in Figure 10: in contrast to other tokens, its cosine similarities are not higher with its adjacent tokens. Given that the [CLS] token behaves like a compressed memory token trained using the next sentence prediction, the phenomenon motivates our current work, suggesting that different priors should be given to special tokens for best performance.
@@ 6 Related Work
GIST Mu et al. (2023) trains LLMs with modified attention mechanisms (similar to Figure 8) to compress prompt information into a few gist tokens. AutoCompressor Chevalier et al. (2023) trains an LLM to recursively compress long prompts by combining compressed tokens with new sub-prompts in each iteration, ultimately collecting all compressed tokens to form a compact representation. ICAE Ge et al. (2024) introduces an AE task enabling LLMs to pre-train compression capabilities on large-scale corpora, requiring only minimal parameter tuning for the encoder while freezing the decoder. 500xCompressor Li et al. (2024b) builds upon ICAE by changing the information carrier from memory token outputs to memory tokens’ KV values. UniICL Gao et al. (2024b) and SelfCP Gao et al. (2024a) freeze both the encoder and decoder and train only a connector module to transform the encoder’s output memory tokens into decoder inputs. VoCo-LLaMA Ye et al. (2025) is the first to use LLMs for compressing visual tokens. Similar to GIST, it compresses visual token information into VoCo tokens through modified attention masks, outperforming methods like Q-Former Li et al. (2023) and average pooling with linear projection Li et al. (2024a). None of these methods discuss the impact of position layout and our EPL can be applied to all these soft prompt methods.
@@
Although we are not aware of position layout work in the compression domain, we find related work in multimodality by considering memory tokens as another modality. Due to images being two-dimensional and text being one-dimensional, handling mixed image-text positional layout remains an open question. Current mainstream approaches in multimodality flatten 2D images into 1D sequences Liu et al. (2023b, a); Bavishi et al. (2023); Lu et al. (2024); Sun et al. (2024). Seeking a more elegant solution, Su (2024) proposes RoPE-Tie by placing visual and text tokens along the diagonal $(y=x)$ in 2D space. Although Su (2024) does not thoroughly validate its design, as the approach maintains the sequence order between modalities and the internal locality of images and text respectively, our empirical results suggest that the design can result in better performance. Qwen2-VL Wang et al. (2024) adopts similar designs in video domains.
@@ 7 Conclusion
We examine position layout, an understudied topic in context compression, and propose EPL for soft prompt methods. EPL improves over default position layout by bringing memory tokens close to its context and at the same time maintains the logical sequence token ordering among context, memory tokens and the subsequent tokens. Extensive experiments show EPL enhances compression efficiency and downstream performance across architectures and modalities.
Given LLM’s widespread use, we believe examining position layout will benefit problems beyond context compression by adopting a prior tailored to tasks that can be expressed in the position layout. We hope EPL’s success fosters research in this underexplored area.
Limitations
Throughout our experiments, we have tested and confirmed EPL’s effectiveness under 5x, 15x, 51x compression ratios for text reconstruction and QA tasks and 4.5x compression ratio for visual QA. There remains a question as to whether our method is still effective at high compression ratio. We don’t have a positive outlook on this question.
First, EPL, requiring memory tokens to achieve uniform coverage across the context, intrinsically aligns with lossless compression scenarios (i.e., autoencoding tasks) which assume that “all tokens in the context are equally important”; however, under high compression ratio (lossy compression), with information carrier capacity being limited, focus should be placed on important tokens, which violates our “equally important” assumption. Empirically, Figure 10 shows that the BERT’s [CLS] would have higher attention to some tokens without a clear pattern. Our compression ratio analysis in Table 4.8 and results in Table 13 on VoCo-LLaMA also show that EPL does not demonstrate any significant gains when used under high compression ratios.
The analysis suggests that high-ratio lossy compression may require a compression mechanism that goes beyond our current work. At a higher level, this suggests that when one adapts position layout methods to different application scenarios, the success can highly depend on whether the conceived layout captures the underlying prior characteristics of specific tasks.
Acknowledgments
This work was supported in part by the National Science Foundation of China (Nos. 62276056 and U24A20334), the Yunnan Fundamental Research Projects (No.202401BC070021), the Yunnan Science and Technology Major Project (No. 202502AD080014), and the Program of Introducing Talents of Discipline to Universities, Plan 111 (No.B16009).
@@ Appendix A Detailed Position Layout
Detailed position layouts are provided in this section. Table 14 illustrates the position layout for LLMs without compression under standard conditions, reflecting natural language priors. Table 15 details the encoder’s position layout. For the decoder, position layouts are described in Tables 16, 17, and 18, with slight variations depending on the specific task (AE, LM, or QA). Each of these tables (Tables 15, 16, 17, and 18) includes a formula at the top representing the generalized layout. Below the formula, an illustrative example is given using a context length $|C|=p=1020$, chunk size $L=510$, compression ratio $r=5$, total sequence length $|X|=2040$, question length $|Q|=50$, and answer length $|A|=5$.
@@ H.1 Sinusoidal Position Encoding
The sinusoidal position encoding Vaswani et al. (2017); Su et al. (2023) is given by: PE_(pos, 2i) = sin(pos100002i/dmodel) PE_(pos, 2i+1) = cos(pos100002i/dmodel)
Consider two nearby positions, pospos and $pos+\delta$, where $\delta$ is a small number.
@@
sin(x + ϵ) ≈sin(x)  for small ϵ cos(x + ϵ) ≈cos(x)  for small ϵ
@@ H.2 Learnable Position Encodings
DPL 0 1 L1L-1 $L$ $L+1$ L+|M|1L+|M|-1
EPL (i1)L+1(i-1)L+1 (i1)L+2(i-1)L+2 iLiL $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor(b+(|M|-1)r\right\rceil$
@@
Table 15: Position Layout of Encoder for $S^{(i)}$. $r=\frac{L}{|M|}$; $b=(i-1)*L+1+\frac{r-1}{2}$. The notation $\left\lfloor\cdot\right\rceil$ indicates rounding to the nearest integer.
@@
DPL(ICAE) 0 1 k|M|1k|M|-1 $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+p$
DPL(500x) $L$ $L+1$ L+|M|1L+|M|-1 $k|M|$ $k|M|+1$ $k|M|+2$ $k|M|+p$
EPL $\left\lfloor b\right\rceil$ $\left\lfloor b+r\right\rceil$ $\left\lfloor b+(k|M|-1)r\right\rceil$ 0 11 22 $p$
@@
EPL 33 88 $1018$ 0 11 22 $1020$
Table 16: Position Layout of Decoder in AE Task. $r=\frac{L}{|M|}$; $b=1+\frac{r-1}{2}$. The notation $\left\lfloor\cdot\right\rceil$ indicates rounding to the nearest integer.
@@
EPL 33 88 $1018$ $1020$ $1021$ $1022$ $2040$
Table 17: Position Layout of Decoder in LM Task. $r=\frac{L}{|M|}$; $b=1+\frac{r-1}{2}$. The notation $\left\lfloor\cdot\right\rceil$ indicates rounding to the nearest integer.
@@
EPL 33 88 $1018$ $1020$ $1021$ $1070$ $1071$ $1075$
Table 18: Position Layout of Decoder in QA Task. $r=\frac{L}{|M|}$; $b=1+\frac{r-1}{2}$; $t=k|M|+|Q|$; $t^{\prime}=|C|+|Q|$. The notation $\left\lfloor\cdot\right\rceil$ indicates rounding to the nearest integer.
@@
Figure 6: Attention Matrix in the 2nd layer of ICAE after MRQA fine-tuning.